Ethanol, used in alcoholic beverages, can be produced by fermentation of sucrose. The balanced equation for the fermentation process is shown below. What mass of ethanol (C2H5OH) would be produced when 2.20 g sucrose reacts by this process? C12H22O11(s) + H2O(l) → 4 C2H5OH(l) + 4 CO2(g)
Answer #1

The molar mass of ethanol = 12(2) + 6 (1) + 16 =46 g/mol.

The molar mass of sucrose = 12(12) + 22(1) + 11 (16) =342 g/mol.

Number of moles are obtained by dividing mass with molar mass.

2.20 g sucrose corresponds to \dfrac {2.20}{342}=0.006433 moles.

1 mole of sucrose gives 4 moles of ethanol.

0.006433 moles of sucrose will give 4 \times 0.006433 = 0.0257 moles of ethanol.

The mass of ethanol produced is obtained by multiplying mass with molar mass

The mass of ethanol produced = 0.0257 \times 46 = 1.18 g